# Steiner Traveling Salesman Problem with Time Windows and Pickup–Delivery: integrating classical and quantum optimization

### v2

@@ 6.3 Model Dimensionality and Computational Complexity
This section analyzes the dimensional growth and computational complexity of the proposed formulations, both in their original form and after applying the AFGR procedure. The analysis quantifies the evolution of the model size in terms of binary and continuous variables, as well as equality and inequality constraints, for increasing problem instances. The results highlight how AFGR effectively reduces model dimensionality, thus improving scalability and enabling the solution of larger instances. To assess the computational viability of the proposed approaches, we conduct an analysis of the two mathematical formulations introduced. The objective of this analysis is to compare the structural characteristics of the models in terms of the number of variables and constraints, which directly influence the scalability and solvability of the problem. The analysis is carried out both (ABF and NBF) with and without the application of the AFGR method, in order to highlight its impact on the structural size of the models. This study is performed across a range of instance sizes, allowing us to evaluate how the problem complexity evolves with increasing numbers of nodes and arcs.
Table 5: Detailed comparison of ABF complexity with and without AFGR method. For each instance size, we report: Bin: number of binary decision variables, Cont: number of continuous variables, Tot Vars: total number of variables (binary + continuous), $Eq:$ number of equality constraints, $\leq$: number of inequality constraints, $\geq$: number of inequality constraints, Tot C: total number of constraints. The columns labeled “With AFGR” show the reduced model size after applying the AFGR. The final columns indicate the percentage reduction in total variables and constraints.
$V$ Without AFGR With AFGR
% Red.
Vars
% Red.
Constr.
4 77 64 141 31 108 84 223 77 64 141 31 108 84 223 0% 0%
5 150 110 260 51 184 160 395 150 110 260 51 184 160 395 0% 0%
6 280 180 460 84 313 308 705 280 180 460 84 313 308 705 0% 0%
7 494 280 774 132 500 551 1183 494 280 774 132 500 551 1183 0% 0%
8 825 416 1241 197 732 900 1829 494 280 774 132 500 551 1183 38% 35%
9 1312 594 1906 284 1064 1440 2788 494 280 774 133 519 570 1222 59% 56%
10 2000 820 2820 395 1489 2200 4084 825 416 1241 199 782 950 1931 56% 53%
11 2940 1100 4040 532 1970 3185 5687 825 416 1241 199 782 950 1931 69% 66%
12 4189 1440 5629 700 2607 4543 7850 1312 594 1906 286 1128 1504 2918 66% 63%
13 5810 1846 7656 901 3372 6300 10573 2000 820 2820 397 1569 2280 4246 63% 60%
14 7872 2324 10196 1138 4277 8528 13943 2940 1100 4040 535 2117 3332 5984 60% 57%
15 10450 2880 13330 1413 5239 11210 17862 2940 1100 4040 535 2117 3332 5984 70% 66%
16 13625 3520 17145 1731 6446 14606 22783 4189 1440 5629 703 2784 4720 8207 67% 64%
17 17484 4250 21734 2094 7828 18724 28646 5810 1846 7656 904 3582 6510 10996 65% 62%
18 22120 5076 27196 2504 9257 23520 35281 7872 2324 10196 1140 4441 8692 14273 63% 60%
19 27632 6004 33636 2966 11008 29359 43333 10450 2880 13330 1416 5524 11495 18435 60% 57%
20 34125 7040 41165 3482 12969 36225 52676 13625 3520 17145 1734 6773 14933 23440 58% 56%
Average Reduction: 47% 44%
Table 6: Detailed comparison of NBF complexity with and without AFGR method. For each instance size, we report: Bin: number of binary decision variables, Cont: number of continuous variables, Tot Vars: total number of variables (binary + continuous), $Eq:$ number of equality constraints, $\leq$: number of inequality constraints, $\geq$: number of inequality constraints, Tot C: total number of constraints. The columns labeled “With AFGR” show the reduced model size after applying the AFGR. The final columns indicate the percentage reduction in total variables and constraints.
$V$ Without AFGR With AFGR
% Red.
Vars
% Red.
Constr.
4 77 64 141 29 108 84 221 77 64 141 29 108 84 221 0% 0%
5 150 110 260 48 184 160 392 150 110 260 48 184 160 392 0% 0%
6 280 180 460 79 313 308 700 280 180 460 79 313 308 700 0% 0%
7 494 280 774 124 500 551 1175 494 280 774 124 500 551 1175 0% 0%
8 825 416 1241 185 732 900 1817 494 280 774 124 500 551 1175 38% 35%
9 1312 594 1906 267 1064 1440 2771 494 280 774 125 519 570 1214 59% 56%
10 2000 820 2820 372 1489 2200 4061 825 416 1241 187 782 950 1919 56% 53%
11 2940 1100 4040 502 1970 3185 5657 825 416 1241 187 782 950 1919 69% 66%
12 4189 1440 5629 662 2607 4543 7812 1312 594 1906 269 1128 1504 2901 66% 63%
13 5810 1846 7656 854 3372 6300 10526 2000 820 2820 374 1569 2280 4223 63% 60%
14 7872 2324 10196 1081 4277 8528 13886 2940 1100 4040 505 2117 3332 5954 60% 57%
15 10450 2880 13330 1345 5239 11210 17794 2940 1100 4040 505 2117 3332 5954 70% 66%
16 13625 3520 17145 1651 6446 14606 22703 4189 1440 5629 665 2784 4720 8169 67% 64%
17 17484 4250 21734 2001 7828 18724 28553 5810 1846 7656 857 3582 6510 10949 65% 62%
18 22120 5076 27196 2397 9257 23520 35174 7872 2324 10196 1083 4441 8692 14216 63% 60%
19 27632 6004 33636 2844 11008 29359 43211 10450 2880 13330 1348 5524 11495 18367 60% 57%
20 34125 7040 41165 3344 12969 36225 52538 13625 3520 17145 1654 6773 14933 23360 58% 56%
Average Reduction: 47% 44%
Table 5 and Table 6 provide a detailed comparison of model complexity for increasing problem sizes ($V$), with and without the application of the AFGR method. The complexity is measured in terms of the total number of binary and continuous variables, as well as equality and inequality constraints.
Without the AFGR method, the size of both models increases rapidly as $|V|$ grows. Both the number of variables and constraints follow a super-linear, nearly exponential trend. For example, when $V$ increases from 4 to 20, the total number of variables rises from 141 to over 41000, and the total number of constraints grows from 223 to more than 52000. This growth illustrates the scalability challenges of the baseline model formulations, especially for large instances.
In contrast, applying the AFGR method significantly slows the growth in model size. While the number of variables and constraints still increases with $V$, the rate is considerably reduced. At $V=20$, for example, AFGR reduces the total number of variables from 41165 to 17145 and constraints from 52676 to 23440 in the ABF model. This corresponds to a 58% reduction in variables and a 56% reduction in constraints. Similar trends are observed in the NBF formulation, with consistent reductions beyond $V\geq 9$.
The benefits of AFGR are particularly evident for larger instance sizes:
  • 1.
    Variables: the original models exhibit near-exponential growth in the number of variables, whereas the AFGR-enhanced versions grow much more gradually, especially beyond $V=10$.
  • 2.
    Constraints: this is where AFGR delivers the most substantial impact, cutting constraint counts by more than half in many cases.
On average, AFGR achieves a 47% reduction in total variables and a 44% reduction in total constraints across all tested values of $V$. This confirms that AFGR not only reduces model size but also significantly improves scalability, enabling the solution of larger problem instances with greater computational efficiency.
(a) Total number of variables in the optimization model with and without AFGR as a function of $|V|$.
(b) Total number of constraints in the optimization model with and without AFGR as a function of $|V|$.
Figure 2: Comparison of the total number of variables (a) and constraints (b) in the optimization models with and without AFGR as a function of $|V|$. Orange lines represent the model with AFGR, while blue lines represent the model without AFGR. Since the ABF and NBF formulations exhibit identical trends in both variables and constraints (as shown in Table 5 and Table 6), the results are reported here in a single representative figure for clarity and conciseness.
Figure 2 illustrate the significant impact of incorporating AFGR on the model size. A consistent reduction is observed in both the total number of variables and constraints across all values of $V$ for both formulations. These results indicate that AFGR enables a more compact model formulation, which is potentially more computationally efficient.
@@ 6.4 Numerical results
The analysis of the results is organized into two main phases, each focusing on a distinct computational setting and on specific structural aspects of the problem.
@@
In each of the above settings, results are reported for both ABF and NBF models, with and without the AFGR procedure, to evaluate the performance impact of the proposed AFGR method. The outcomes include solution quality, constraint satisfaction and runtime analysis. Comparative tables and figures are provided to support the discussion and highlight the advantages and trade-offs among the different formulations and problem variants.
@@ 6.4.1 Performance analysis with a classical solver
We present a comprehensive computational analysis structured in four main stages. First, we examine the performance of the proposed formulations on the complete STSP-TWPD, assessing their scalability and the impact of the AFGR preprocessing method. Second and third, we extend the analysis to the two simplified variants, STSP-PD and STSP-TW, which are derived from the main formulation by selectively removing time-window or pickup-and-delivery constraints, respectively. Finally, we provide a comparative discussion among the three models (STSP-TWPD, STSP-PD, and STSP-TW), highlighting how their structural and computational characteristics evolve under different constraint configurations.
@@
We evaluate the computational performance of the STSP-TWPD model with two formulations (ABF and NBF). Table 7 reports the computational results obtained for the ABF and NBF formulations across instances of increasing size $V$. For each formulation, we present the objective value (OF), the computational time, and the optimality gap returned by Gurobi under a time limit of 5000 seconds.
Table 7: Comparison between the ABF and the NBF on instances of increasing size $V$. For each formulation, we report: OF, Time and GAP Opt. All experiments are performed with Gurobi using a time limit of 5000 seconds.
$V$ ABF NBF
OF Time (s) GAP Opt OF Time (s) GAP Opt
4 965.69 0.00 0% 965.69 0.02 0%
5 995.96 0.06 0% 995.96 0.03 0%
6 1019.62 0.08 0% 1019.62 0.09 0%
7 1401.16 3.13 0% 1401.16 3.52 0%
8 1429.73 6.84 0% 1429.73 5.89 0%
9 1299.21 12.05 0% 1299.21 5.97 0%
10 1620.41 742.86 0% 1620.41 177.25 0%
11 1559.06 2079.08 0% 1559.06 1019.55 0%
12 1324.50 136.00 0% 1324.50 68.70 0%
13 1062.23 49.47 0% 1062.23 31.92 0%
14 1350.57 3046.52 0% 1350.57 8207.86 0%
15 1333.68 2140.61 0% 1333.68 2279.67 0%
16 1498.74 15767.30 19% 1498.74 18817.78 16%
The results in Table 7 show that, for all instances where optimality is proven within the time limit, the ABF and NBF formulations return identical objective values. This empirical evidence supports the theoretical equivalence previously established in subsection 4.3.1. With respect to computational time, the two formulations display comparable behavior, with differences that are not systematic and can be attributed to solver-dependent branching decisions.
For instances with $V\geq 10$, both models exhibit a marked increase in computational effort, reflecting the growing complexity of the STSP–TWPD. The variability in running times across instances also suggests that performance is influenced not only by model size but by specific structural characteristics of each test case.
The instance with $V=16$ is particularly informative: although both formulations identify the same best integer solution, neither is able to prove optimality within the allotted time, resulting in nonzero optimality gaps. The difference between the reported gaps (19% for ABF and 16% for NBF) does not reflect discrepancies in solution quality but instead arises from the distinct tightness of their respective linear relaxation bounds. In other words, while the integer solutions coincide, the relaxations produce different lower bounds, which directly translate into the observed differences in optimality gaps.
We also assess the contribution of the AFGR method in improving the tractability and efficiency of the solution process. The experiments are conducted on instances with up to $V=10$ customers. Each model is tested in two configurations: with and without AFGR. All models are solved using Gurobi, with a time limit of 5 seconds for each instance. Table 8 report the results for the ABF and NBF formulations.
Table 8: ABF and NBF formulations of the STSP-TWPD model with and without the AFGR module. Each row corresponds to an instance with $V$ customers. For each formulation (ABF and NBF), results are shown both with and without AFGR. Metrics include: Vars: number of decision variables, FO: average objective function value across ten runs, Time [s]: average solution time in seconds, GAP opt: average optimality gap with respect to the best-known solution. The last two columns report the percentage variable reduction (% var red) for ABF and NBF.
$V$ ABF NBF
Without AFGR With AFGR % var red Without AFGR With AFGR % var red
Vars FO Time [s] GAP opt Vars FO Time [s] GAP opt Vars FO Time [s] GAP opt Vars FO Time [s] GAP opt
4 145 965,69 0,02 0% 145 965,69 0,00 0% 0% 145 965,69 0,02 0% 145 965,69 0,00 0% 0%
5 265 995,96 0,02 0% 265 995,96 0,02 0% 0% 265 995,96 0,00 0% 265 995,96 0,00 0% 0%
6 466 1019,62 0,06 0% 466 1019,62 0,08 0% 0% 466 1019,62 0,03 0% 466 1019,62 0,05 0% 0%
7 781 1401,16 2,95 0% 781 1401,16 3,78 0% 0% 781 1401,16 2,97 0% 781 1401,16 3,00 0% 0%
8 1249 1429,73 4,58 0% 781 1429,73 2,13 0% 37% 1249 1429,73 5,13 0% 841 1429,73 2,78 0% 33%
9 1915 1299,21 6,33 0% 781 1299,21 1,64 0% 59% 1915 1299,21 5,45 0% 901 1299,21 1,48 0% 53%
10 2830 1639,03 19,47 28% 1249 1620,41 16,84 0% 56% 2830 1678,47 19,41 43% 1405 1620,41 11,89 0% 50%
It is important to clarify that the reported optimality gap (GAP opt) values correspond to the Mixed-Integer Programming (MIP) gap returned by the classic solver Gurobi at termination. The MIP gap is a standard performance metric in mixed-integer optimization, defined as the relative difference between the best feasible solution (BestObj) and the best known lower bound (BestBound) at the time the solver stops:
$\text{MIP Gap (\%)}=100\times\frac{|\text{BestObj}-\text{BestBound}|}{|\text{BestObj}|}$ (48)
This metric quantifies how close the current solution is to proven optimality. A non-zero MIP gap typically indicates that the solver reached the predefined time limit before closing the gap entirely, rather than that the model provides a weak relaxation.
Both ABF and NBF formulations are able to find optimal solutions for instances up to $V=9$, with zero optimality gap. For $V=10$, both formulations show a marked increase in computational complexity. In the full ABF model (without AFGR), the solver returns a solution with a 28% optimality gap. The corresponding full NBF model performs worse, reaching a 43% gap. These results highlight the inherent difficulty of the problem at larger sizes and the need for strategies to improve tractability. Introducing AFGR significantly improves the performance of both formulations at larger instance sizes. In the ABF model, AFGR reduces the variable count by 56% at $V=10$, and enables the solver to close the optimality gap completely, achieving a 0% gap in the same time limit. A similar trend is observed in the NBF model, where AFGR achieves a 50% reduction in variables and reduces the optimality gap from 43% to 0%. In addition to improving solution quality, AFGR also reduces the average computational time in larger instances. For example, the average runtime for ABF at $V=10$ drops from 19.47s (without AFGR) to 16.84s (with AFGR), while for NBF it drops from 19.41s to 11.89s. This suggests that the simplified search space generated by AFGR improves solver efficiency without compromising solution quality. AFGR introduces no changes in small instances ($V\leq 7$), but leads to substantial reductions in the variables size as $V$ increases. The reduction reaches up to 56% for ABF and 50% for NBF at $V=10$, illustrating the effectiveness of AFGR in pruning infeasible or redundant decision variables in dense scenarios.
@@
We evaluate the performance of the STSP-PD using both the ABF and NBF formulations, tested with and without the AFGR module. Table 9 reports the detailed results for instances with up to $V=10$ customers.
Table 9: ABF and NBF formulations of the STSP-PD with and without the AFGR. Each row corresponds to an instance with $V$ customers. For each formulation (ABF and NBF), results are shown both with and without AFGR. Metrics include: Vars: number of decision variables, FO: average objective function value across ten runs, Time [s]: average solution time in seconds, GAP opt: average optimality gap with respect to the best-known solution. The last two columns report the percentage variable reduction (% var red) for ABF and NBF.
$V$ ABF NBF
Without AFGR With AFGR % var red Without AFGR With AFGR % var red
Vars FO Time [s] GAP opt Vars FO Time [s] GAP opt Vars FO Time [s] GAP opt Vars FO Time [s] GAP opt
4 145 965,69 0,00 0% 145 965,69 0,05 0% 0% 145 965,69 0,03 0% 145 965,69 0,05 0% 0%
5 265 995,96 0,00 0% 265 995,96 0,02 0% 0% 265 995,96 0,00 0% 265 995,96 0,02 0% 0%
6 466 1019,62 0,08 0% 466 1019,62 0,13 0% 0% 466 1019,62 0,13 0% 466 1019,62 0,09 0% 0%
7 781 1011,19 0,09 0% 781 1011,19 0,11 0% 0% 781 1011,19 0,08 0% 781 1011,19 0,05 0% 0%
8 1249 980,07 0,20 0% 781 980,07 0,13 0% 37% 1249 980,07 0,23 0% 841 980,07 0,20 0% 33%
9 1915 937,29 0,17 0% 781 937,29 0,09 0% 59% 1915 937,29 0,27 0% 901 937,29 0,20 0% 53%
10 2830 941,05 1,52 0% 1249 941,05 0,53 0% 56% 2830 941,05 1,48 0% 1405 941,05 0,81 0% 50%
All configurations solve to optimality within a negligible amount of time, with no impact from AFGR on either solution time or variable count. This confirms that, in the absence of time windows, the problem becomes significantly easier to solve, and both formulations remain equivalent in terms of performance and solution quality under these conditions.
As the number of customers increases, AFGR begins to significantly reduce the number of decision variables, up to 56% for ABF and 50% for NBF at $V=10$. Despite this reduction, all instances still reach optimality with a 0% gap, highlighting that removing time windows greatly simplifies the solution space. Moreover, AFGR contributes to lower solution times for larger instances. For instance, in the ABF formulation at $V=10$, the average time drops from 1.52s to 0.53s with AFGR; similarly, NBF drops from 1.48s to 0.81s. These results confirm that, even though the model complexity grows with $V$, the absence of temporal constraints makes the problem highly manageable from a computational standpoint.
When comparing these results with those obtained from the complete model including time windows (Table 8), the impact of the additional constraints becomes evident, particularly for instances with $V\geq 8$. In the complete model, solution times increase significantly, and the solver fails to close the optimality gap within the time limit for $V=10$ in both formulations without AFGR: the ABF model reaches a 28% gap, while the NBF model reaches 43%. In contrast, in the no-time-windows model, both formulations still achieve a 0% gap even at $V=10$, clearly indicating that time windows introduce considerable combinatorial complexity. AFGR proves to be a valuable component in both settings. However, its impact is more pronounced when time windows are included: in the complete model, applying AFGR restores optimality at $V=10$, while also halving the number of variables in both formulations. Nevertheless, solution times remain significantly higher compared to the no-time-windows scenario. For example, the ABF model with AFGR takes 16,84s at $V=10$, compared to 0,53s in the no-time-windows case.
Thus, time windows are an important factor in computational difficulty: their removal drastically reduces both solution times and model size. AFGR contributes meaningfully in both cases, but its utility is particularly evident when time windows are present, as it helps recover tractability in otherwise hard instances. The objective values remain identical across the two scenarios for each $V$, confirming that time windows do not alter the cost structure, but rather constrain the feasible set and increase solver effort.

### v3

@@ 6.3 Model Dimensionality and Computational Complexity
We provide a comprehensive analysis of the dimensional growth and computational complexity of the proposed formulations, ABF and NBF, for the STSP-TWPD, both in their original form and after the application of the AFGR reduction method. The objective of this analysis is twofold: first, to quantify how the structural size of the models evolves as a function of instance characteristics; second, to rigorously assess the effectiveness of AFGR in controlling such growth across heterogeneous classes of instances.
The analysis is conducted over all instance classes defined in Table 4, and the results are reported in A (Tables 615). These tables cover a wide spectrum of structural configurations, ranging from sparse to highly constrained networks. This setting allows us to jointly evaluate the impact of instance size, expressed in terms of the number of nodes $|V|$, while model dimensionality is measured in terms of binary and continuous variables, as well as equality and inequality constraints, thus providing a complete and consistent characterization of each formulation. This comprehensive evaluation framework ensures that the observed trends are robust across heterogeneous structural conditions.
The results in Tables 615 clearly indicate that, across all instance classes, the dimensionality of both formulations scales rapidly with the problem size when AFGR is not applied. For the ABF formulation, increasing the number of nodes from $|V|=4$ to $|V|=20$ leads to a sharp growth in the number of variables (from 145 to 41,185) and constraints (from 258 to 80,853). The largest values are consistently observed in denser configurations such as A1, B1, and B2. This is explained by the higher arc density in these instances, which directly amplifies the number of arc–time combinations and, consequently, the number of routing variables and associated constraints. This confirms that arc density plays a critical role in driving model dimensionality, beyond the mere increase in the number of nodes.
A fully analogous behavior is observed for the NBF formulation, confirming that both models share the same structural growth pattern. Although ABF and NBF exhibit the same theoretical order of growth, they differ in their practical dimensionality. In particular, ABF tends to generate a larger number of constraints due to the explicit arc-based indexing, whereas NBF achieves a slightly more compact constraint representation while preserving the same asymptotic complexity. This suggests that, while the two formulations are theoretically equivalent, their structural organization has a non-negligible impact on practical model size.
The effect of AFGR is both substantial and systematic across all instances. Indeed, the reduction ranges from a minimum of approximately 30% in the number of variables and 31% in the number of constraints for instances such as A1 and B1, up to peaks of 71% (variables) and 73% (constraints) for instances A5 and B5.
A more detailed analysis across instance classes reveals a clear and systematic dependency of the reduction effectiveness on the structural properties of the graph. In sparse instances (e.g., A1–A2 and B1–B2), where the ratio $|V_{r}|/|V|$ is relatively low, the reduction is moderate, typically between 30% and 44%. In these cases, a larger portion of the network must be preserved to ensure connectivity among required nodes. This suggests that the reduction potential is inherently limited when the structure of the graph is already close to the minimal connectivity requirements. In intermediate configurations (e.g., A3 and B3), the reduction becomes more pronounced, generally around 55%. Finally, in dense and highly constrained instances (e.g., A4–A5 and B4–B5), the reduction reaches its highest levels, approaching and in some cases exceeding 70%. It is also important to note that these latter classes correspond to the largest instances ($|V|=20$), where the reduction becomes even more dramatic, reaching up to approximately 90% in the number of variables and 91% in the number of constraints. This provides compelling evidence that AFGR becomes increasingly effective as both instance size and structural density grow.
This behavior is a direct consequence of the AFGR mechanism. By retaining only arcs belonging to shortest paths between required nodes, AFGR systematically removes structurally redundant connections. As the proportion of required nodes increases, the relative number of irrelevant arcs grows significantly, allowing the reduction process to become increasingly selective and effective. This suggests that AFGR implicitly exploits structural redundancy in dense graphs, leading to a more aggressive pruning of the arc set. As a result, the arc set is drastically reduced, leading to a substantial decrease in both variables and constraints.
The detailed growth patterns are further illustrated in B (Figures 312), which provide a visual representation of the evolution of model size across all instance classes. In all cases, the original formulations exhibit a steep and clearly increasing trend in both variables and constraints, with a marked acceleration for $|V|\geq 12$. In contrast, the curves corresponding to the AFGR-enhanced models remain significantly flatter, indicating a much more controlled scaling behavior. This visual evidence reinforces the quantitative results observed in A. A particularly relevant observation is that the gap between the original and reduced models widens progressively as $|V|$ increases. While the difference is relatively limited for small instances, it becomes substantial for larger ones (e.g., $|V|\geq 15$), where the reduced models require only a fraction of the variables and constraints of the original formulations. This clearly indicates that the benefits of AFGR scale positively with problem size, making it particularly suitable for large-scale instances. This effect is especially evident in Figures 6, 7, 11, and 12, corresponding to instances A4, A5, B4, and B5, respectively, where the reduction in both variables and constraints is particularly pronounced. This confirms that AFGR is especially effective in dense and highly constrained scenarios, where structural redundancy is more significant.
@@ 6.4 Numerical results
The analysis of the results is organized into two main phases.
@@
In each of the above settings, results are reported for both ABF and NBF models, with and without the AFGR procedure, to evaluate the performance impact of the proposed AFGR method.
@@ 6.4.1 Performance analysis with a classical solver
We present a comprehensive computational analysis structured in four main stages. First, we examine the performance of the proposed formulations on the complete STSP-TWPD, assessing their scalability and the impact of the AFGR preprocessing method. Second and third, we extend the analysis to the two simplified variants, STSP-PD and STSP-TW, respectively. Finally, we provide a comparative discussion among the three models (STSP-TWPD, STSP-PD, and STSP-TW), highlighting how their structural and computational characteristics evolve under different constraint configurations.
The computational performance of the three proposed models is evaluated using both the ABF and NBF formulations across all instance classes reported in Table 4. A time limit of 600 seconds is imposed for all experiments.
@@
The complete set of results is reported in C.1 (Tables 1625).
A first fundamental observation emerging from the computational analysis is that, for the vast majority of tested instances, both ABF and NBF formulations with and without AFGR, achieve the same objective function value. This result confirms the theoretical equivalence of the two formulations demonstrated in Section 4.3.1 and, more importantly, shows that the application of the AFGR procedure does not deteriorate solution quality.
A detailed analysis reveals an exception in the most challenging instances. In particular, for instance class $A2$, the increasing problem size leads to a progressive deterioration in the performance of the standard formulations. While for smaller instances both ABF and NBF are able to reach optimality, significant differences emerge as $|V|$ increases. For larger instances ($|V|\geq 14$), the standard formulations begin to exhibit non-zero optimality gaps, indicating a deterioration in solution quality.
For example, at $|V|=14$, the standard NBF formulation reaches an objective value of 1352,80 with a gap of 11%, whereas the reduced model attains the optimal value of 1350,57 with zero gap. This discrepancy becomes more pronounced for larger instances. At $|V|=16$, the standard formulations show gaps up to 27–28%, with objective values of 1551,41, while the reduced models significantly improve the solution quality, reaching values as low as 1498,74 with zero gap.
A similar trend is observed for $|V|=17$: the standard formulations exhibit optimality gaps of approximately 23–24%, with an objective value of 1295,56, whereas the AFGR-based formulations attain optimal solutions (gap equal to 0%) and a slightly improved objective value of 1294,31. An even more significant improvement is observed at $|V|=18$, where the standard formulations report gaps up to 34%, while the reduced models converge to the optimal value of 1250,49 with zero gap. At $|V|=19$, the standard formulations show gaps up to 35%, while AFGR reduces both the objective value and the gap (e.g., for ABF formulation, from 1317,96 to 1308,23, with gap reduction from 35% to 6%).
Finally, for $|V|=20$, both standard formulations (ABF and NBF) attain an objective value of 1380.82. By contrast, the application of AFGR leads to a significant improvement, reducing the objective value to 1358.66 for ABF (23% reduction) and to 1348.40 for NBF (19% reduction).
A similar behaviour can be observed for instance class $B2$. For instance, at $|V|=14$, the standard NBF formulation attains an objective value of 1352,80 with a gap of 9%, whereas the reduced model achieves the optimal value of 1350,57 with zero gap. This behaviour becomes more pronounced as the instance size increases.
At $|V|=16$, both standard formulations exhibit a gap of 27%, with an objective value of 1551,41, while the reduced models reach a significantly better solution of 1498,74 with zero gap. A similar trend is observed for $|V|=17$ and $|V|=18$. Focusing on the ABF formulation, the standard models exhibit optimality gaps of up to 15% and 34%, respectively, whereas the AFGR-based models consistently achieve optimal solutions (zero gap) together with improved objective values (e.g., from 1295.56 to 1294.31 for $|V|=17$, and from 1251.55 to 1250.49 for $|V|=18$). The advantage remains evident for the largest instances. At $|V|=19$, the standard formulations report gaps up to 28%, whereas the reduced models either achieve optimality or significantly reduce the gap.
Finally, for $|V|=20$, the standard formulations exhibit optimality gaps as high as 37–40%. In contrast, the application of AFGR leads to substantially improved solutions, reducing both the optimality gap (down to 18–12%) and the objective value (from 1376,16 to 1358,66 for ABF, and from 1376,16 to 1348,40 for NBF).
These results clearly demonstrate that the reduced models obtained through AFGR consistently achieve solutions with zero or significantly lower optimality gaps, often improving upon the solutions found by the standard formulations.
From a computational perspective, the impact of AFGR on solution times is particularly significant. The reduction procedure leads to substantial improvements in computational efficiency for both formulations, with time savings that become increasingly pronounced as the problem size grows. More specifically, the average time reduction achieved by AFGR ranges between approximately 52% and 98%, depending on the instance class and formulation. For smaller instances (e.g., class $A1$ and $B1$), the average time saving is already considerable, reaching about 61% for ABF and 66% for NBF. As the problem size increases, the benefits of AFGR become even more evident. For medium-sized instances (classes A2A2B3B3), the reduction exceeds 52% and reaches up to 95%. Finally, for the largest instances (classes A4A4B5B5), the computational gain becomes dramatic, with average time savings of up to 98% for both formulations.
Overall, these results clearly highlight the effectiveness of the AFGR preprocessing method. While preserving optimality across all tested instances, it drastically reduces computational effort, making the solution of large-scale STSP-TWPD instances significantly more tractable. This confirms that preprocessing techniques such as AFGR are not only beneficial but essential when dealing with complex routing problems characterized by time windows and pickup-and-delivery constraints.
@@
This part of the paper presents the computational results for the STSP-PD variant, with particular focus on the comparison between standard and AFGR-based formulations. We also analyze the impact of removing time-window constraints on model performance and solution quality. The full set of results is provided in C.2 (Tables 2635).
A first key observation is that both formulations consistently achieve identical objective function values across all tested instances. This behavior can be attributed to the reduced complexity of the problem setting, as the absence of time window constraints makes the model significantly easier to solve compared to the full STSP-TWPD variant.
From a computational standpoint, the impact of AFGR on solution times is substantial. The preprocessing procedure yields very significant reductions in computational time for both formulations, with improvements that increase as the instance size grows. For the smallest instances (classes $A1$ and $B1$), the average time savings are already notable, reaching approximately 68% for ABF and 77% for NBF. As the instance size increases (classes A2A2B3B3), the computational gains become even more pronounced, with reductions ranging between 85% and over 90%. In particular, for class $A3$ and $B3$, the time savings reach up to 93% for the NBF formulation. The most remarkable improvements are observed in the largest instances (classes A4A4B5B5), where the reduction becomes dramatic. In these cases, AFGR allows for average time savings between 93% and 97% for both formulations, effectively reducing solution times by more than an order of magnitude. This highlights the crucial role of preprocessing in enhancing scalability for larger and more complex instances.
A comparison of the average objective function values between STSP-TWPD and STSP-PD reveals a clear impact of time window constraints on solution quality. In particular, for several instance classes (e.g., A2 and B2), the STSP-TWPD exhibits significantly higher objective values than STSP-PD, reflecting the additional routing restrictions imposed by time windows. Moreover, while STSP-PD consistently yields identical objective values across all formulations and configurations, STSP-TWPD shows slight variations between standard and reduced models, with AFGR providing marginal improvements.
A further comparison between the STSP-TWPD and STSP-PD variants highlights a substantial difference in computational effort. As expected, the presence of time-window constraints in STSP-TWPD leads to significantly higher solution times compared to the STSP-PD model. This gap is particularly evident in medium-sized instances. For example, in class $A2$, the average solution time for the standard formulation reaches approximately 982 seconds for ABF and over 1000 seconds for NBF in STSP-TWPD, whereas the corresponding values for STSP-PD remain below 50 seconds. A similar behavior is observed for class $B2$, where the STSP-TWPD requires more than 1000 seconds for all instance sizes, while the STSP-PD remains around 50 seconds. This clearly indicates that the inclusion of time-window constraints dramatically increases the computational complexity of the problem. Even for instances classes $A3$ and $B3$, the difference remains significant. STSP-TWPD requires between 118 and 129 seconds in the standard formulation for all instance sizes, while STSP-PD solves the same instances in approximately half the time (around 50–78 seconds). For instances classes $A1$, $B1$, A4A4B5B5, the gap is less extreme but still consistent. In all cases, STSP-TWPD systematically requires more computational time than STSP-PD, confirming the additional burden introduced by temporal constraints. The impact of AFGR is remarkable in both variants. For STSP-TWPD, the preprocessing reduces solution times from hundreds or even thousands of seconds to single-digit values in many cases (e.g., from over 1000 seconds to less than 10 seconds in classes $A2$ and $B2$). A similar effect is observed for STSP-PD, although starting from much lower initial times.
Thus, as expected, the STSP-TWPD variant generally requires higher computational times in its standard formulation, due to the additional complexity introduced by time-window constraints. However, despite this increased modeling complexity, the AFGR procedure effectively mitigates the computational burden, yielding performance gains that are fully comparable to those observed in the STSP-PD case.